At the beginning of the year, a high school football coach decided to leave his job and give up his annual coaching salary of $55,000 and open his own sporting goods store. A partial income statement for follows: 3. Revenues Revenue from sales of goods and services Operating costs and expenses: .... $210,000 Cost of products and services sold Selling expenses Administrative expenses Total operating costs and expenses Income from operations Interest expense (bank loan) Non-recurring expenses to start business. $82,000 $6,000 $12,000 S100,000 si 10,000 $14,000 $24,000 $72,000 To get the sporting goods store opened, the former coach used $60,000 of his personal savings. The coach opened his store in a building that he owns. Prior to opening his store, the building was rented for $36,000 per year. The coach could have earned 5 percent return by investing in stocks of other new businesses with risk levels similar to the risk level associated with his new sporting goods store. a. The former high school coach incurs S of total explicit costs for using market-supplied resources. b. The opportunity cost of the owners equity capital is S c. Total implicit cost of owner-supplied resources is S d. Total economic cost is S 1 By how much did coachs wealth change by opening the sporting resources is Sannually , and accounting profit is S
Answer #1

Answer- 3:

a) Explicit costs are out of pocket costs for a firm. Former high school coac incur following costs for marked supplied resources:

Cost of Product and Services Sold

82000

Selling Expense

6000

Administrative Expense

12000

Interest Expense

14000

Non-recurring expense to start business

24000

Total Explicit costs paid

138000

b) Opportunity Cost of Owner’s Equity Capital:

Opportunity cost is value of the next best alternative foregone when a different alternative is taken. As the owner can invest his capital in stocks with same risk level with 5% return, so the opportunity cost of capital is:

= 60,000 * 5% = $ 3,000 annually.

c) Implicit Cost of Owner Supplied Resources:

Implicit costs are the opportunity cost of resources owned by owner and used in business i.e. for which no cash payment is made

Opportunity Cost of Capital (60,000 * 5%)

3000

Opportunity Cost Building used as Store (Rent)

36000

Opportunity Cost of high school football coach (salary)

55000

Total Implicit Costs

94000

       

?

d) Total Economic Cost and Accounting Profit

Economic Cost = Explicit Costs + Implicit Costs

= 138,000 + 94,000 = $ 232,000

Accounting Profit = Income from Operations – Interest Expense

= 110,000 – 14,000 = $ 96,000

Note: Non-recurring expense are capitalised and written off over a long period.

e) Change in Wealth of Coach:

The coach has earned a profit of $ 96,000 by starting his own sporting goods store. The opportunity cost of opening the store is $ 94,000 i.e. benefits foregone as result of opening store. Hence change in wealth:

= 96,000 – 94,000 = $ 2,000.